# Optimizing Posterior Samples for Bayesian Optimization via Rootfinding

### v2

@@ 3.1 TS-roots Algorithm
$\widetilde{f}(\mathbf{x};\widetilde{\omega})=f(\mathbf{x};\omega)+b(\mathbf{x};\widetilde{\omega}),\quad f(\mathbf{x};\omega)=\prod_{i=1}^{d}f_{i}(x_{i};\omega_{i}),\quad b(\mathbf{x};\widetilde{\omega})=\sum_{j=1}^{n}v_{j}\kappa(\mathbf{x},\mathbf{x}^{j}).$ (2)
Here, the prior sample $f(\mathbf{x})$ is a separable function determined by the random bits $\omega$. Data adjustment $b(\mathbf{x})$ is a sum of the canonical basis with coefficients $\mathbf{v}=(\mathbf{K}_{n,n}+\boldsymbol{\Sigma})^{-1}(\mathbf{y}-\mathbf{f}_{n}-\boldsymbol{\varepsilon})$. Both the data adjustment and the posterior sample are determined by the random bits $\widetilde{\omega}=(\omega,\boldsymbol{\varepsilon})$. In the following, we denote the prior and the posterior samples as $f_{\omega}$ and $\widetilde{f}_{\widetilde{\omega}}$, respectively. Our goal is to find the global minimum $(\widetilde{\mathbf{x}}_{\widetilde{\omega}}^{\star},\widetilde{f}_{\widetilde{\omega}}^{\star})$ of the posterior sample $\widetilde{f}_{\widetilde{\omega}}(\mathbf{x})$.
@@ 6 Results
In each optimization iteration, we record the best observed value of the error $\log(y_{\min}-f^{\star})$ and the distance $\log\left(\left\|\mathbf{x}_{\min}-\mathbf{x}^{\star}\right\|\right)$, where $y_{\min}$, $\mathbf{x}_{\min}$, $f^{\star}$, and $\mathbf{x}^{\star}$ are the best observation of the objective function in each iteration, the corresponding location of the observation, the true minimum value of the objective function, and the true minimum location, respectively. We compare the optimization results obtained from TS-roots and other BO methods, including GP-TS using decoupling sampling with random Fourier features (TS-DSRF), GP-TS with random Fourier features (TS-RF), expected improvement (EI) [1], and lower confidence bound (LCB)—the version of GP-UCB [5] for minimization problems.
Figure 2 shows the medians and interquartile ranges of solution values obtained from 20 runs of each of the considered BO methods. The corresponding histories of solution locations are in Figure 9 of Appendix I. With a fair comparison of outer-loop results (detailed in Appendix H), TS-roots surprisingly shows the best performance on the 2D Schwefel, 16D Ackley, and 16D Powell functions, and gives competitive results in the 4D Rosenbrock and 10D Levy problems. Notably, TS-roots recommends better solutions than its counterparts, TS-DSRF and TS-RF, in high-dimensional problems and offers competitive performance in low-dimensional problems. Across all the examples, EI and LCB tend to perform better in the initial stages, while TS-roots shows fast improvement in later stages. This is because GP-TS favors exploration, which delays rewards. The exploration phase, in general, takes longer for higher-dimensional problems.
@@
Figure 2(f) shows the outer-loop optimization results for the truss obtained from 20 runs of each BO method, where $f^{\star}$ is a lower bound of the best objective function value we observed from all runs. TS-roots provides the best optimization result with rapid convergence.
@@ C.1 Spectral Sampling of Separable Gaussian Process Priors
Given spectral representations of the univariate component functions of a separable Gaussian Process prior, we can accurately approximate the prior sample as:
$f_{\omega}(\mathbf{x})=\prod_{i=1}^{d}f_{i}(x_{i};\omega_{i}),\quad f_{i}(x_{i};\omega_{i})\approx\sum_{k=0}^{N_{i}-1}w_{i,k}\sqrt{\lambda_{i,k}}\phi_{i,k}(x_{i}).$ (10)
Here $N_{i}$ is selected for each variate such that $\lambda_{i,N_{i}-1}/\lambda_{i,1}\leq\eta_{i}$, where $\eta_{i}$ is sufficiently small (see Appendix H for the value used in this study). Using spectral representations of the univariate components as in Eq. 10 is much more efficient than directly using a spectral representation of the separable GP prior, because the former uses $\sum_{i=1}^{d}N_{i}$ univariate terms to exactly represent $\prod_{i=1}^{d}N_{i}$ multivariate terms in the latter.
@@ Appendix G Ten-bar Truss
Consider a ten-bar truss shown in Figure 8. The truss has ten members and is subjected to vertical load $P_{1}=60$ kN at node 2, vertical load $P_{2}=40$ kN at node 3, and horizontal load $P_{3}=40$ kN at node 3. The Young’s modulus of the truss material $E=200$ GPa. The length parameter $L=1$ m. Let $A({\bf x})=\sum_{i=1}^{10}x_{i}$ and $\delta_{3}({\bf x})$ denote the total area of the cross-sectional areas of the truss members and the vertical displacement at node 3, respectively, where ${\bf x}=[x_{1},\dots,x_{10}]^{\intercal}$ is the vector of cross-sectional areas of the truss members. The optimization problem formulated for the truss is to minimize both $A({\bf x})$ and $\delta_{3}({\bf x})$. Since $A({\bf x})$ and $\delta_{3}({\bf x})$ are competing, we define the objective function as a weight-sum of $A({\bf x})$ and $\delta_{3}({\bf x})$, such that
@@ Appendix I Additional Results
Figures 10 and 11 compare the performance of the inner-loop optimization by three different initialization schemes, i.e., rootfinding, uniform grid, and Latin hypercube sampling, for low-dimensional cases of the 2D Schwefel and 4D Rosenbrock functions, and for higher-dimensional cases of the 10D Levy, 16D Ackley, and 16D Powell functions. Rootfinding performs better than the uniform grid and Latin hypercube sampling initialization schemes, especially in high-dimensional settings.

### v4

@@ 3.1 TS-roots Algorithm
$\widetilde{f}(\mathbf{x};\widetilde{\omega})=f(\mathbf{x};\omega)+b(\mathbf{x};\widetilde{\omega}),\quad f(\mathbf{x};\omega)\approx\prod_{i=1}^{d}f_{i}(x_{i};\omega_{i}),\quad b(\mathbf{x};\widetilde{\omega})=\sum_{j=1}^{n}v_{j}\kappa(\mathbf{x},\mathbf{x}^{j}).$ (2)
Here, the prior sample $f(\mathbf{x})$ is approximated as a separable function determined by the random bits $\omega$ (see Section C.1). Data adjustment $b(\mathbf{x})$ is a sum of the canonical basis with coefficients $\mathbf{v}=(\mathbf{K}_{n,n}+\boldsymbol{\Sigma})^{-1}(\mathbf{y}-\mathbf{f}_{n}-\boldsymbol{\varepsilon})$. Both the data adjustment and the posterior sample are determined by the random bits $\widetilde{\omega}=(\omega,\boldsymbol{\varepsilon})$. In the following, we denote the prior and the posterior samples as $f_{\omega}$ and $\widetilde{f}_{\widetilde{\omega}}$, respectively. Our goal is to find the global minimum $(\widetilde{\mathbf{x}}_{\widetilde{\omega}}^{\star},\widetilde{f}_{\widetilde{\omega}}^{\star})$ of the posterior sample $\widetilde{f}_{\widetilde{\omega}}(\mathbf{x})$.
@@ 6 Results
In each optimization iteration, we record the best observed value of the error $\log(y_{\min}-f^{\star})$ and the distance $\log\left(\left\|\mathbf{x}_{\min}-\mathbf{x}^{\star}\right\|\right)$, where $y_{\min}$, $\mathbf{x}_{\min}$, $f^{\star}$, and $\mathbf{x}^{\star}$ are the best observation of the objective function in each iteration, the corresponding location of the observation, the true minimum value of the objective function, and the true minimum location, respectively. We compare the optimization results obtained from TS-roots and other BO methods, including GP-TS using decoupling sampling with random Fourier features (TS-DSRF), GP-TS with random Fourier features (TS-RF), expected improvement (EI) [1], lower confidence bound (LCB)—the version of GP-UCB [5] for minimization problems, and LogEI [43].
Figure 2 shows the medians and interquartile ranges of solution values obtained from 20 runs of each of the considered BO methods. The corresponding histories of solution locations are in Figure 10 of Appendix I. With a fair comparison of outer-loop results (detailed in Appendix H), TS-roots surprisingly shows the best performance on the 2D Schwefel, 16D Ackley, and 16D Powell functions, and gives competitive results in the 4D Rosenbrock and 10D Levy problems. Notably, TS-roots recommends better solutions than its counterparts, TS-DSRF and TS-RF, in high-dimensional problems and offers competitive performance in low-dimensional problems. Across all the examples, EI and LCB tend to perform better in the initial stages, while TS-roots shows fast improvement in later stages. This is because GP-TS favors exploration, which delays rewards. The exploration phase, in general, takes longer for higher-dimensional problems. TS-roots outperforms LogEI on the 2D Schwefel, 4D Rosenbrock, 10D Levy, and 16D Ackley functions (detailed in Figure 9 of Appendix I).
@@
Figures 2(f) and 9(f) show the outer-loop optimization results for the truss obtained from 20 runs of each BO method, where $f^{\star}$ is a lower bound of the best objective function value we observed from all runs. TS-roots provides the best optimization result with rapid convergence.
@@ 7 Conclusion and Future Work
Acknowledgments
We would like to thank Tian PAN for his help with the maxk_sum problem, and the organizers and the participants of the NeurIPS 2024 Workshop on Bayesian Decision-making and Uncertainty for their invaluable discussions and feedback on an earlier version of the paper. The authors are supported by the University of Houston through the SEED program no. 000189862.
@@ C.1 Spectral Sampling of Separable Gaussian Process Priors
Given spectral representations of the univariate covariance functions of a separable Gaussian Process prior, we can accurately approximate the prior sample as:
$f_{\omega}(\mathbf{x})\approx\prod_{i=1}^{d}f_{i}(x_{i};\omega_{i}),\quad f_{i}(x_{i};\omega_{i})\approx\sum_{k=0}^{N_{i}-1}w_{i,k}\sqrt{\lambda_{i,k}}\phi_{i,k}(x_{i}).$ (10)
Here $N_{i}$ is selected for each variable such that $\lambda_{i,N_{i}-1}/\lambda_{i,1}\leq\eta_{i}$, where $\eta_{i}$ is sufficiently small (see Appendix H for the value used in this study). The product of the univariate sample functions forms a third-order approximation of the multivariate sample function; their moment functions differ at order four and higher even orders. Using spectral representations of the univariate components as in Eq. 10 is much more efficient than directly using a spectral representation of the separable GP prior, because the former uses $\sum_{i=1}^{d}N_{i}$ univariate terms to approximate $\prod_{i=1}^{d}N_{i}$ multivariate terms in the latter. Another benefit of this representation is that the sample functions are separable, which can be exploited in downstream tasks.
@@ Appendix G Ten-bar Truss
Consider a ten-bar truss shown in Figure 8. The truss has ten members and is subjected to vertical load $P_{1}=60$ kN at node 2, vertical load $P_{2}=40$ kN at node 3, and horizontal load $P_{3}=10$ kN at node 3. The Young’s modulus of the truss material $E=200$ GPa. The length parameter $L=1$ m. Let $A({\bf x})=\sum_{i=1}^{10}x_{i}$ and $\delta_{3}({\bf x})$ denote the total area of the cross-sectional areas of the truss members and the vertical displacement at node 3, respectively, where ${\bf x}=[x_{1},\dots,x_{10}]^{\intercal}$ is the vector of cross-sectional areas of the truss members. The optimization problem formulated for the truss is to minimize both $A({\bf x})$ and $\delta_{3}({\bf x})$. Since $A({\bf x})$ and $\delta_{3}({\bf x})$ are competing, we define the objective function as a weight-sum of $A({\bf x})$ and $\delta_{3}({\bf x})$, such that
@@ Appendix I Additional Results
Comparison of TS-roots and LogEI Outer-loop Optimization Results.
Figure 9 compares the performance of the outer-loop optimization of TS-roots and LogEI [43] on the five benchmark functions and a real-world ten-bar truss structure. Across all examples, save 16D Powell, TS-roots outperforms LogEI, further demonstrating the robustness of our approach. The comparatively weaker performance on the 16D Powell function can be attributed to its convex landscape, where BO generally underperforms relative to gradient-based methods.
@@
Figures 11 and 12 compare the performance of the inner-loop optimization by three different initialization schemes, i.e., rootfinding, uniform grid, and Latin hypercube sampling, for low-dimensional cases of the 2D Schwefel and 4D Rosenbrock functions, and for higher-dimensional cases of the 10D Levy, 16D Ackley, and 16D Powell functions. Rootfinding performs better than the uniform grid and Latin hypercube sampling initialization schemes, especially in high-dimensional settings.
@@
Figure 9: Outer-loop optimization results for the (a) 2D Schwefel function, (b) 4D Rosenbrock function, (c) 10D Levy function, (d) 16D Ackley function, (e) 16D Powell function, and (f) ten-bar truss problem. The plots are histories of medians and interquartile ranges of solution values from 20 runs of TS-roots and LogEI.